# Learning Latent Space Dynamics with Model-Form Uncertainties: A Stochastic Reduced-Order Modeling Approach

### v1

@@ 1 Introduction
Constructing efficient and robust surrogate models for decision-making and engineering system design is a critical task in computational science and engineering. In this paper, we outline a generally applicable approach for representing and propagating model-form uncertainties by leveraging principles and tools from reduced-order modeling, Riemannian optimization and information theory. This type of uncertainties arises from the inherent limitations of approximating complex physical phenomena using mathematical models. They originate from simplifying assumptions [31], lack of knowledge [30, 11, 23], and/or specification of a training strategy [47, 16], among others. By construction, such uncertainties cannot be captured in a parametric setting [43] and require ad hoc stochastic representations that encode the action of the underlying learned model. The methodology presented in this paper extracts, from data, knowledge and intuition that is not captured by a conventional, deterministic computational model and incorporates it into the modeling approach in a non-intrusive fashion.
@@
The original formulation of the OpInf methodology calls for a proper orthogonal decomposition (POD) to be applied to the data matrix to yield an orthonormal basis that spans a subspace of fixed dimensionality [37]. The POD provides an optimal linear embedding of our original high-dimensional data set into a reduced-dimensional linear subspace [42, 6]. These methods thus face significant challenges when applied to problems in which the Kolmogorov N-width decreases slowly with increasing subspace dimension [13]. The Kolmogorov-width is a mathematical measure characterizing the system’s reducibility through its best linear approximation error. Limitations surrounding the use of linear subspaces become particularly problematic in systems with strong nonlinearities, moving discontinuities, or parametric dependencies. To address this issue, nonlinear MOR methods that utilize machine learning methods have been successfully applied to several applications in the field of reduced-order modeling. Despite the success of neural-network-based nonlinear MOR techniques, these methods often require significant computational resources and sizable amount of data for training [33, 40, 7, 14]. In contrast, various recent works [3, 19, 17, 18] have studied a nonlinear manifold-based approach that enrich state approximations with low-order polynomial terms, leading to improved accuracy and stability of ROMs. We will follow the principles behind the latter in the development of our probabilistic approach.
@@
Despite recent efforts aimed at establishing probabilistic versions of data-driven ROMs by endowing the underlying reduced-order operators with uncertainty, significant gaps remain. Perhaps the most natural step towards the integration of uncertainties consists in casting the calibration task as a Bayesian inverse problem; see [27] for example, as well as [34] for a Bayesian formulation leveraging enrichment through Gaussian processes. The work by Uy et al. [50] has addressed operator inference with noisy data. Progress has also been made with regard to probabilistic error estimates; see [49] for linear OpInf ROMs in the case reduced-order operators are equivalent to those obtained from explicit projection-based methods as in [36]. Note that providing an exhaustive review on UQ methods in the more general context of scientific machine learning (SciML), or for other classes of operator approximations (using, e.g., deep learning), is beyond the scope of this work; see [38] for a general review articulating SciML and UQ, [51, 15, 57] for deep-learning formulations embedding UQ aspects, and [9] for Bayesian nonlocal operator regression, to list a few.
This work considers the nonlinear manifold-based Operator Inference (OpInf) framework proposed in [18]. Building upon recent developments pertaining to stochastic reduced-order modeling for nonlinear dynamical systems [44, 45, 54] and coupled partial differential equations [53], we develop a methodological extension to the OpInf formulation that combines the randomization of a projection matrix, via Riemannian projection and retraction operators, with an information-theoretic probabilistic model. The proposed method enables the sampling of model perturbations with proper concentration in the convex hull defined by a set of model proposals, identified by variations in the selection of training parameters for instance. This strategy offers several key advantages:
@@
By addressing the representation of model-form uncertainties in OpInf methods, this work contributes to the broader goal of enhancing the robustness of learning approaches for reduced-order models in complex physics-based systems. An overview of the proposed probabilistic framework is shown in Fig. 1.
@@ 2.1 Learning Nonlinear Representations
Consider a training dataset comprised of a set of snapshots with dimension $N$ corresponding to the high-dimensional state-space variable (e.g. temperature or pressure) in a physical system of interest. We denote each snapshot by $\boldsymbol{s}_{j}\in\mathbb{R}^{N}$, and assemble a snapshot matrix $[S]\in\mathbb{R}^{N\times k}$ from $k$ such snapshots:
@@
Note that $[S]$ may contain states from multiple full model evaluations corresponding to different initial conditions and/or boundary conditions. To enable the derivation of the probabilistic formulation in a rather general setting, we consider low-dimensional approximation $\Gamma:\mathbb{R}^{r}\mapsto\mathbb{R}^{N}$ to ${\boldsymbol{s}}(t)\in\mathbb{R}^{N}$ defined as
@@
where $\hat{\boldsymbol{s}}(t)\in\mathbb{R}^{r}$ is the reduced-order state vector at time $t\in\mathbb{R}$, ${\boldsymbol{s}}_{\text{ref}}$ is a reference state vector (defined through, e.g., an initial condition or the mean system-state), $[V_{r}]\in\mathbb{R}^{N\times r}$ and $[\overline{V}]\in\mathbb{R}^{N\times q}$ are such that $[[V_{r}],[\overline{V}]]\in\mathbb{R}^{N\times(r+q)}$ is a projection matrix computed from the snapshot matrix $[S]$, with $(r+q)\ll N$ typically, and $\mathcal{N}:\mathbb{R}^{r}\mapsto\mathbb{R}^{q}$ defines a nonlinear enrichment for the state-space approximation. Several choices for $\mathcal{N}$ have been proposed in the literature (see Remark 1 at the end of this section). Here, we consider the polynomial approximation defined in [18] as
@@
where the coefficient matrix $[\Xi]\in\mathbb{R}^{q\times(p-1)r}$ weights the contributions of the singular vectors in $[\overline{V}]$ and the vector $\boldsymbol{g}(\hat{\boldsymbol{s}}(t))\in\mathbb{R}^{(p-1)r}$ has components $\hat{\boldsymbol{s}}^{j}(t)=[\hat{\boldsymbol{s}}_{1}^{j}(t),\hat{\boldsymbol{s}}_{2}^{j}(t),\dots,\hat{\boldsymbol{s}}_{r}^{j}(t)]^{\top}$ where $j\in[2,\dots,p]$. The snapshots of the reduced-order state-space variable $\hat{\boldsymbol{s}}(t)$ are obtained by solving the following constrained optimization problem
@@
and $\text{St}(N,(r+q))$ is the Stiefel manifold. The $\gamma$-regularization term in Eq. (4) is introduced to avoid overfitting. The representation learning problem (4)–(5) can be solved using a number of techniques (see [17, 18], for instance).
It should be noted that linear approximations given by projections onto the leading $r$ principal components can lead to poor results in combination with correction terms because the linear subspace is agnostic to the nonlinear feature map. This problem may be alleviated by selecting principal components that are not necessarily ordered in descending order with respect to the singular values. The greedy method introduced in [41] demonstrates that such a strategy has the potential to increase the representation accuracy by several orders of magnitude without compromising on scalability or computational efficiency. This work, however, employs the first $r$ leading principal components of the training data for building the linear part of the approximation. This is done to isolate and analyze the effects of uncertainties in the reduced basis construction.
@@ 2.2 Learning Latent Space Dynamics
where $\mathbf{f}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ maps the system state to its time derivative and $\boldsymbol{s}_{0}\in\mathbb{R}^{N}$ is a given initial condition. In the general case where $\mathbf{f}(\boldsymbol{s})$ is non-polynomial, variable transformations may be introduced that promote a system structure that is more amenable to projection-based approximation. This procedure is often referred to as lifting [5, 32, 39, 21]. It has been demonstrated in [26] that many engineering systems are characterized by nonlinear terms that may be lifted to linear-quadratic form:
@@
where $\otimes$ denotes the Kronecker product. By incorporating nonlinear state approximation (2) into the above equation and applying a Galerkin projection, an approximation of the latent dynamics with polynomial structure is obtained:
@@
where $\hat{\boldsymbol{c}}\in\mathbb{R}^{r}$, $[\hat{A}]\in\mathbb{R}^{r\times r}$, $[\hat{H}]\in\mathbb{R}^{r\times r^{2}}$, and $[\hat{P}]\in\mathbb{R}^{r\times d(r,p)}$ define reduced-order matrix operators. $\hat{\boldsymbol{s}}(0)=\hat{\boldsymbol{s}}_{0}$ is the representation of initial condition $\boldsymbol{s}_{0}$ in the latent space. The vector $\hat{\boldsymbol{g}}(\hat{\boldsymbol{s}})$ is a subvector of $\hat{\boldsymbol{s}}$ (of length $d(r,p)$) and contains monomials of degree three up to degree 2p. The matrix $[\hat{P}]$ accounts for higher-order interactions between the reduced-order state coefficients, captured by the vector $\hat{\boldsymbol{g}}(\hat{\boldsymbol{s}})$. Details regarding its construction can be found in [18].
@@
The nonnegative scalars $\{\lambda_{i}\}_{i=1}^{3}$ are Tikhonov regularization parameters that are introduced to promote stability in the inferred operator and prevent overfitting [35]. The time derivative of the reduced state vector is estimated numerically using finite differences.
@@ 3.1.1 Preliminaries
The matrix $[B]^{\top}$ represents a boundary condition operator that enforces physical constraints on the system. These constraints are inherent to the original full-order model and must be preserved in the reduced-order representation. Specifically, the full-order solution $s$ is assumed to satisfy
@@
In order to test the robustness of the learning framework with respect to the training strategy, we consider $m$ training scenarios giving raise to a family $\mathcal{A}$ of $m$ projection matrices, $\mathcal{A}=\{[\Phi^{(1)}],\dots,[\Phi^{(m)}]\}$, termed the anchor points in the following. For later use, we introduce the following sets:
@@
with $[V_{r}^{(i)}]=[V_{r}(\mu^{(i)})]$ and $[\overline{V}^{(i)}]=[\overline{V}(\mu^{(i)})]$. We also introduce the global projection matrix $[\Phi^{*}]=[[V_{r}^{*}]\ [\overline{V^{*}}]]\in\mathbb{S}_{N,(r+q)}$, the definition of which is discussed in Section 3.2.
@@ 3.1.2 Probabilistic Representation on a Subset of the Stiefel Manifold
Following the standard modeling approach on $\text{St}(N,(r+q))$ (e.g., for interpolation with parametric partial differential equations), the probabilistic formulation is developed using the tangent space to the manifold at basepoint $[\Phi^{*}]$. This construction necessitates the choice of projection and retraction operators, written as
@@
respectively. While many choices were proposed for these operators in the literature [1], it was shown in [54] that the use of iterative algebraic approximations to the Riemannian projection and retractions operators, denoted by $\exp_{[\Phi^{*}]}^{\text{St}}$ (Riemannian exponential) and $\log_{[\Phi^{*}]}^{\text{St}}$ (Riemannian logarithm) respectively, preserves the action of the linear constraints in the following sense:
@@
    If $[M]\in\mathbb{S}_{N,(r+q)}$, then $P_{[\Phi^{*}]}([M])=\exp_{[\Phi^{*}]}^{\text{St}}([M])\in\mathbb{S}_{N,(r+q)}$;
@@
    If $[\Delta]\in\mathbb{S}_{N,(r+q)}$, then $R_{[\Phi^{*}]}([\Delta])=\log_{[\Phi^{*}]}^{\text{St}}([\Delta])\in\mathbb{S}_{N,(r+q)}$;
where we recall that $[\Phi^{*}]\in\mathbb{S}_{N,(r+q)}$ by construction. Based on this observation, and using the fact that the tangent space is a vector space, it follows that
@@
Eq. (19) provides a path towards an admissible randomization of the projection matrix $[\Phi]=[[V_{r}][\overline{V}]]$. Before proceeding with the definition of the underlying probability measure, it is instructive to indicate some advantages and disadvantages associated with the use of Riemannian operators. Starting with the latter, it should be noticed that the computation of such operators can be intensive for large values of $N$, $r$ and $q$. While the use of the approximation algorithms in [56] facilitate deployment for scenarios of practical interest, further work is necessary to accelerate the sampling task for the above representation. On the other hand, the use of these operators offer two important advantages. First, they facilitate statistical inference and in particular, the prescription of the Fréchet mean for the fluctuations on the manifold (see Section 3.2). Second, imposing that the coefficients of $\mathbf{p}$ define a convex combination on the tangent space $T_{[\Phi^{*}]}\text{St}(N,(r+q))$, it can be shown (under assumptions defined in [2] that the random version of Eq. (19) yields samples located in the convex hull defined by the anchor points—which provides a simple interpretation of the probabilistic model, regardless of the dimensions.
@@
The underlying probability measure $\pi$ is derived in the context of information theory, using the principle of maximum entropy under repulsive constraints defining the convex combination on the tangent space. Specifically, the consideration of the constraints
@@
where $\mathbb{E}$ is the operator of mathematical expectation, leads to the consideration of a Dirichlet distribution with concentration parameters collected in $\boldsymbol{\alpha}$, $\pi=\mathcal{D}(\boldsymbol{\alpha})$.

### v2

@@ 1 Introduction
Constructing efficient and robust surrogate models for decision-making and engineering system design is a critical task in computational science and engineering. In this paper, we outline a generally applicable approach for representing and propagating model-form uncertainties by leveraging principles and tools from reduced-order modeling, Riemannian optimization and information theory. This type of uncertainty arises from the inherent limitations of approximating complex physical phenomena using mathematical models. They originate from simplifying assumptions [33], lack of knowledge [32, 13, 25], and/or specification of a training strategy [54, 18], among others. By construction, such uncertainties cannot be captured in a parametric setting [50] and require ad hoc stochastic representations that encode the action of the underlying learned model. The methodology presented in this paper extracts, from data, knowledge and intuition that is not captured by a conventional, deterministic computational model and incorporates it into the modeling approach in a non-intrusive fashion.
@@
The original formulation of the OpInf methodology calls for a proper orthogonal decomposition (POD) to be applied to the data matrix to yield an orthonormal basis that spans a subspace of fixed dimensionality [42]. The POD provides an optimal linear embedding of our original high-dimensional data set into a reduced-dimensional linear subspace [49, 6]. These methods thus face significant challenges when applied to problems in which the Kolmogorov N-width decreases slowly with increasing subspace dimension [15]. The Kolmogorov-width is a mathematical measure characterizing the system’s reducibility through its best linear approximation error. Limitations surrounding the use of linear subspaces become particularly problematic in systems with strong nonlinearities, moving discontinuities, or parametric dependencies. To address this issue, nonlinear MOR methods that utilize machine learning methods have been successfully applied to several applications in the field of reduced-order modeling. Despite the success of neural-network-based nonlinear MOR techniques, these methods often require significant computational resources and sizable amount of data for training [38, 45, 8, 16]. In contrast, various recent works [3, 21, 19, 20] have studied a nonlinear manifold-based approach that enriches state approximations with low-order polynomial terms, leading to improved accuracy and stability of ROMs. We will follow the principles behind the latter in the development of our probabilistic approach.
@@
Several contributions aiming at establishing probabilistic versions of data-driven ROMs (to accommodate noisy data, for instance) have been proposed recently. Perhaps the most natural step towards the integration of uncertainties consists in casting the calibration task as a Bayesian inverse problem; see [29] for example, as well as [39] for a Bayesian formulation leveraging enrichment through Gaussian processes. The work by Uy et al. [57] has addressed operator inference with noisy data. Progress has also been made with regard to probabilistic error estimates; see [56] for linear OpInf ROMs in the case reduced-order operators are equivalent to those obtained from explicit projection-based methods, as in [41]. Note that providing an exhaustive review on UQ methods in the more general context of scientific machine learning (SciML), or for other classes of operator approximations (using, e.g., deep learning), is beyond the scope of this work; see [43] for a general review articulating SciML and UQ, [58, 17, 64] for deep-learning formulations embedding UQ aspects, and [11] for Bayesian nonlocal operator regression, to list a few.
This work considers the nonlinear manifold-based Operator Inference (OpInf) framework proposed in [20] and seeks to enhance its robustness by allowing for the probabilistic description of uncertainties in predictions as the training strategy is being refined, e.g., as the parametric combinations indexing the training dataset vary in both values and cardinality. Since each combination defines a given reduced-order operator, we interpret the above uncertainties as model-form uncertainties and thus aim to construct a probabilistic model in a nonparametric setting. Building upon recent developments pertaining to stochastic reduced-order modeling for nonlinear dynamical systems [51, 52, 61] and coupled partial differential equations [60], we specifically develop a methodological extension to the OpInf formulation that combines the randomization of the projection matrix, which effectively expands the approximation space, with a probabilistic model defined in the context of information theory [34, 35]. In what follows, this model will be referred to as the information-theoretic model. The proposed method enables the sampling of model perturbations with proper concentration in the convex hull defined by a set of model proposals, identified by variations in the selection of training parameters for instance. This strategy offers several key advantages:
@@
By addressing the representation of model-form uncertainties in OpInf methods, this work contributes to the broader goal of enhancing the robustness of learning approaches for reduced-order models in complex physics-based systems. In doing so, it complements other contributions focusing on model error correction (see, e.g., [7, 36]) and optimal architecture selection (see [48] for a contribution in the context of Bayesian model selection, for instance), developed for deep learning models. An overview of the proposed probabilistic framework is shown in Fig. 1.
@@ 2.1 Learning Nonlinear Representations
Consider a training dataset comprised of a set of snapshots corresponding to the high-dimensional state-space variable (e.g. temperature or pressure) in a physical system of interest. We denote each snapshot by $\boldsymbol{s}_{j}\in\mathbb{R}^{N}$, and assemble a snapshot matrix $[S]\in\mathbb{R}^{N\times k}$ from $k$ such snapshots:
@@
Note that $[S]$ may contain states from multiple full model evaluations corresponding to different initial conditions and/or boundary conditions. To enable the derivation of the probabilistic formulation in a rather general setting, we consider a low-dimensional approximation $\Gamma:\mathbb{R}^{r}\mapsto\mathbb{R}^{N}$ to ${\boldsymbol{s}}(t)\in\mathbb{R}^{N}$ defined as
@@
where $\hat{\boldsymbol{s}}(t)\in\mathbb{R}^{r}$ is the reduced-order state vector at time $t$, defined as the coefficients of the reduced model in the low-dimensional (linear subspace) basis, ${\boldsymbol{s}}_{\text{ref}}$ is a reference state vector (defined through, e.g., an initial condition or the mean system-state), $[V_{r}]\in\mathbb{R}^{N\times r}$ and $[\overline{V}]\in\mathbb{R}^{N\times q}$ are orthonormal matrices such that $[[V_{r}],[\overline{V}]]\in\mathbb{R}^{N\times(r+q)}$ is a projection matrix computed from the snapshot matrix $[S]$, with $(r+q)\ll N$ typically, and $\mathcal{N}:\mathbb{R}^{r}\mapsto\mathbb{R}^{q}$ defines a nonlinear enrichment for the state-space approximation.
Several choices for the enrichment function $\mathcal{N}$ have been proposed in the literature (see Remark 1 at the end of this section). Here, we consider the polynomial approximation defined in [20] as
@@
where the coefficient matrix $[\Xi]\in\mathbb{R}^{q\times(p-1)r}$ weights the contributions of the singular vectors in $[\overline{V}]$ and the vector $\boldsymbol{g}(\hat{\boldsymbol{s}}(t))\in\mathbb{R}^{(p-1)r}$ has components $\hat{\boldsymbol{s}}^{j}(t)=[\hat{\boldsymbol{s}}_{1}^{j}(t),\hat{\boldsymbol{s}}_{2}^{j}(t),\dots,\hat{\boldsymbol{s}}_{r}^{j}(t)]^{\top}$, with $j\in[2,\dots,p]$. The snapshots of the reduced-order state-space variable $\hat{\boldsymbol{s}}(t)$ are obtained by solving the following constrained optimization problem
@@
$\gamma$ is a regularization parameter, and $\text{St}(N,(r+q))$ is the Stiefel manifold (which is the set of $N$-by-$(r+q)$ orthonormal matrices; see the reference paper [10], as well as [1], for details). The $\gamma$-regularization term in Eq. (4) is introduced to avoid overfitting [20]. The representation learning problem (4)–(5) can be solved using a number of techniques (see [19, 20], for instance).
It should be noted that linear approximations given by projections onto the leading $r$ principal components may lead to poor results in combination with correction terms because the linear subspace is agnostic to the nonlinear feature map. This problem may be alleviated by selecting principal components that are not necessarily ordered in descending order with respect to the singular values. The greedy method introduced in [46] demonstrates that such a strategy has the potential to increase the representation accuracy by several orders of magnitude without compromising on scalability or computational efficiency. This work, however, employs the first $r$ leading principal components of the training data for building the linear part of the approximation. This is done to isolate and analyze the effects of uncertainties in the reduced basis construction.
@@ 2.2 Learning Latent Space Dynamics
where $\mathbf{f}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ maps the system state $\boldsymbol{s}$ (with values in $\mathbb{R}^{N}$) to its time derivative and $\boldsymbol{s}_{0}\in\mathbb{R}^{N}$ is a given initial condition. In the general case where $\mathbf{f}$ is non-polynomial, variable transformations may be introduced that promote a system structure that is more amenable to projection-based approximation. This procedure is often referred to as lifting and was described in [5, 37, 44, 23]. It has been demonstrated in [28] that many engineering systems are characterized by nonlinear terms that may be lifted to a linear-quadratic form:
@@
where $\otimes$ denotes the Kronecker product [^1] The Kronecker product is defined as an operation on two matrices of arbitrary size. In our case, the matrix $\boldsymbol{s}$ consists of a single column representing the system state. Acknowledging the slight abuse of notation, the Kronecker product $\boldsymbol{s}\otimes\boldsymbol{s}$ can be viewed as vectorization of the outer product., $[A]\in\mathbb{R}^{N\times N}$, and $[H]\in\mathbb{R}^{N\times N^{2}}$. By incorporating the nonlinear state approximation (2) into the above equation, and applying a Galerkin projection, an approximation of the latent dynamics with polynomial structure is obtained:
@@
where $\hat{\boldsymbol{c}}\in\mathbb{R}^{r}$, $[\hat{A}]\in\mathbb{R}^{r\times r}$, $[\hat{H}]\in\mathbb{R}^{r\times r^{2}}$, and $[\hat{P}]\in\mathbb{R}^{r\times d(r,p)}$ define reduced-order matrices, and $\hat{\boldsymbol{s}}(0)=\hat{\boldsymbol{s}}_{0}$ is the representation of initial condition $\boldsymbol{s}_{0}$ in the latent space. The vector $\hat{\boldsymbol{g}}(\hat{\boldsymbol{s}})$ is a subvector of $\hat{\boldsymbol{s}}$ (of length $d(r,p)$) and contains monomials of degree three up to degree $2p$. The matrix $[\hat{P}]$ accounts for higher-order interactions between the reduced-order state coefficients, captured by the vector $\hat{\boldsymbol{g}}(\hat{\boldsymbol{s}})$. Details regarding its construction can be found in [20].
@@
The nonnegative scalars $\{\lambda_{i}\}_{i=1}^{3}$ are Tikhonov regularization parameters that are introduced to promote stability in the inferred operator and prevent overfitting [40]. This measure enhances the capability of the ROM for issuing reliable predictions outside the regime in which the data were acquired. The time derivative of the reduced state vector can be estimated numerically, using finite differences for instance.
@@ 3.1.1 Preliminaries
The matrix $[B]^{\top}$ represents a boundary condition operator that enforces physical constraints on the system. These constraints are inherent to the original full-order model and must be preserved in the reduced-order representation. Specifically, the full-order solution $\boldsymbol{s}$ is assumed to satisfy
@@
In order to test the robustness of the learning framework with respect to the training strategy, we consider $m$ training scenarios (selected according to the procedure defined in Section 3.1.3) giving rise to a family $\mathcal{A}$ of $m$ projection matrices, $\mathcal{A}=\{[\Phi^{(1)}],\dots,[\Phi^{(m)}]\}$, termed the anchor points in the following. For later use, we introduce the following sets:
@@
with $[V_{r}^{(i)}]=[V_{r}(\mu^{(i)})]$ and $[\overline{V}^{(i)}]=[\overline{V}(\mu^{(i)})]$. We also introduce the global projection matrix $[\Phi^{*}]=[[V_{r}^{*}]\ [\overline{V^{*}}]]\in\mathbb{S}_{N,(r+q)}$, the definition of which is discussed in Section 3.1.3.
@@ 3.1.2 Probabilistic Representation on a Subset of the Stiefel Manifold
Following the standard modeling approach on $\text{St}(N,(r+q))$ (e.g., for interpolation with parametric partial differential equations), the probabilistic formulation is developed using the tangent space $T_{[\Phi^{*}]}\text{St}(N,(r+q))$ to the manifold $\text{St}(N,(r+q))$ at basepoint $[\Phi^{*}]$ (see, e.g., [10] for definition and properties). From a methodological standpoint, this construction first necessitates the choice of projection and retraction operators, written as
@@
respectively (see, e.g., Chapter 5 in [1]). While many choices were proposed for these operators in the literature [1], it was shown in [61] that the use of iterative algebraic approximations to the Riemannian projection and retractions operators, denoted by $\log_{[\Phi^{*}]}^{\text{St}}$ (Riemannian logarithm) and $\exp_{[\Phi^{*}]}^{\text{St}}$ (Riemannian exponential) respectively, preserves the action of the linear constraints in the following sense:
@@
    If $[M]\in\mathbb{S}_{N,(r+q)}$, then $P_{[\Phi^{*}]}([M])=\log{[\Phi^{*}]}^{\text{St}}([M])\in\mathbb{S}_{N,(r+q)}$;
@@
    If $[\Delta]\in\mathbb{S}_{N,(r+q)}$, then $R_{[\Phi^{*}]}([\Delta])=\exp{[\Phi^{*}]}^{\text{St}}([\Delta])\in\mathbb{S}_{N,(r+q)}$;
with $[\Phi^{*}]\in\mathbb{S}_{N,(r+q)}$ (as assumed in Section 3.1.1). Based on this observation, and using the fact that the tangent space is a vector space, it follows that
@@
Eq. (19) provides a path towards an admissible randomization of the projection matrix $[\Phi]=[[V_{r}][\overline{V}]]$. Before proceeding with the definition of the underlying probability measure, it is instructive to indicate some advantages and disadvantages associated with the use of Riemannian operators. Starting with the latter, it should be noticed that the computation of such operators can be intensive for large values of $N$, $r$ and $q$. While the use of the approximation algorithms in [63] facilitate deployment for scenarios of practical interest, further work is necessary to accelerate the sampling task for the above representation. On the other hand, the use of these operators offer two important advantages. First, they facilitate statistical inference and in particular, the prescription of the Fréchet mean for the fluctuations on the manifold (see Section 3.2). Second, imposing that the coefficients of $\mathbf{p}$ define a convex combination on the tangent space $T_{[\Phi^{*}]}\text{St}(N,(r+q))$, it can be shown (under assumptions defined in [2]) that the random version of Eq. (19) yields samples located in the convex hull defined by the anchor points—which provides a simple interpretation of the probabilistic model, regardless of the dimensions.
@@
The underlying probability measure $\pi$, which implicitly defines the stochastic model on the Stiefel manifold, is derived in the context of information theory [34, 35], using the principle of maximum entropy [47]. Following the above discussion and leveraging the results derived in [2], we introduce repulsive constraints defining a convex combination on the tangent space:
@@
where $\mathbb{E}$ is the operator of mathematical expectation. These constraints lead to the consideration of a Dirichlet distribution with concentration parameters collected in $\boldsymbol{\alpha}$, $\pi=\mathcal{D}(\boldsymbol{\alpha})$. Note that the probability measure defining the matrix-valued random variable $[\mathbf{\Phi}]$, while not determined explicitly due to the nonlinearity of the mappings involved, is fully determined by the measure $\pi$ on the tangent space, as well as the pullback action of the Riemannian retraction operator.